Circular No. 8981

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 202P/SCOTTI S. Nakano (2004, Nakano Note No. 1070) suggested that P/2001 X2 =

1929 WW, an apparently asteroidal object for which approximate positions from Lowell Observatory plates taken by E. C. Slipher and C. W. Tombaugh on 1929 Nov. 27 and Dec. 3 were published by H. L. Giclas (1940, A.N. 271, 43). Accurate measurements by B. A. Skiff and C. M. Olmstead were published on MPC 16527 in 1990. Although the identification was plausible, the linkage required use of the approximate position on the first night, and this differs from the accurate position by some 6′. Skiff has now reexamined that first plate, finding the measurement on MPC 16527 in fact to apply to a string of faint stars. His new estimate of the position of the object reported by Giclas,

1929 UT  α_{2000}  δ_{2000}  Mag.
Nov. 27.25556  5^{h}17^{m}18.^{s}5  +20^{o}21′06′′  16.5

clearly confirms the correctness of the identification with the object now designated comet 202P (IAUC 8971, 8976), this object having T = 1930 Jan. 8.7 TT. Indeed, Skiff remarks that the object appears weakly cometary on both 1929 plates, the fuzziness amounting to a 10′′ coma at best, with no indication of a tail.

COMETS C/2008 J13–J16, C/2008 K1, C/2008 K2 (SOHO) Further to IAUC 8964, additional Kreutz sungrazers have been found

on SOHO website images. Karl Battams notes that C/2008 J13 was slightly diffuse (mag ∼ 6.5) with a short tail. C/2008 J14 was stellar in appearance (mag ∼ 5.5–6) in C3 images, and was bright and teardrop-shaped in C2 images. C/2008 J15 was small and slightly diffuse (mag ∼ 8). C/2008 J16 was extremely faint (mag ∼ 8.5) and very diffuse. C/2008 K1 was tiny and stellar in appearance (mag ∼ 8–8.5). C/2008 K2 was tiny, extremely faint (mag ∼ 8.5), and slightly diffuse.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 J13 May 12.021 3^{h}20.^{m}2 +16^{o}41′ C2  RK 2008-M11
C/2008 J14  12.971 3 33.9 +15 00 C3/2 RM 2008-M11
C/2008 J15  13.064 3 24.2 +16 54 C2  JR 2008-M11
C/2008 J16  14.379 3 28.6 +17 11 C2  JR 2008-M11
C/2008 K1  17.317 3 39.7 +17 49 C2  MK 2008-M13
C/2008 K2  17.650 3 40.8 +17 52 C2  MK 2008-M13
2008 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green